1. Calculating inflation using a simple price index Consider a fictional price index, the College Student Price Index (CSPI), based on a typical college students annual purchases. Suppose the following table shows information on the market basket for the CSPI and the prices of each of the goods in 2014, 2015, and 2016. The cost of each item in the basket and the total cost of the basket are shown for 2014. Perform these same calculations for 2015 and 2016, and enter the results in the following table 2014 2015 2016 Price Cost Price Cost Price Cost Quantity in Basket (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) Notebooks Calculators Large coffees Energy drinks Textbooks Total cost Price index 15 2 70 2 2 120 30 70 500 100 1,200 1,900 100 5 8 100 130 250 50 10 4 6 150 180

Suppose the base year for this price index is 2014. In the last row of the table, calculate and enter the value of the CSPI for the remaining years Between 2014 and 2015, the CSPI increased by % . Between 2015 and 2016, the CSPI increased by Which of the following, if true, would illustrate why price indexes such as the CSPI might overstate inflation in the cost of going to college? Check all that apply. As the price of calculators rose, fewer students decided to buy them, opting instead to use the free calculators in their cell phones or on their computers. O Professors required each student to buy 10 textbooks, regardless of the price. Energy drinks became increasingly popular on college campuses between 2014 and 2016 due to significant improvements in flavor, but this quality change is hard to measure. O A new mobile device for personal computing became available for purchase.
Answer #1

The following tabular calculation shows the year-on-year basis of basket of goods purchased by the students and related calculation of the inflation rate. It has been found that price of the basket is increasing significantly each year. The inflation rate between 2014 and 2015 is 25%. Inflation rate between 20015 and 2016 is 20%.

2014 2015 2016 Price Cost Price Cost Price Cost Quantity in Basket (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) (Dollars) Notebooks Calculators Large coffees Energy drinks Textbooks Total cost Price index 15 2 30 70 500 100 15*5=75 1*100 100 250-2-500 50*4 200 15*8 = 120 1*130 130 250*2-500 50*6 300 10* 180 1800 2850 8 100 2 4 70 130 250 50 10 2 6 120 1,20o 15010-150-15010*180 -180 2375 1,900 (2375/1900)*100 -125 100 (2850/1900) 100 150

Between 2014 and 2015, CSPI increased by = (125-100)/100 = 25%

Between 2015 and 2016, CSPI increased by = (150-125)/125 = 20%

Correct Answer:

B

C

Textbooks and energy drinks are being purchased as demanded, even if the price is rising significantly. It makes the index to overstate the inflation.